Rationalize
10 divided by the sqrt. of 21 + sqrt of 11
: Rationalize
10 divided by the sqrt. of 21 + sqrt of 11

Answer by HyperBrain(507)

10/(sqrt(21)+sqrt(11))=(10/(sqrt(21)+sqrt(11)))((sqrt(21)-sqrt(11))/(sqrt(21)-sqrt(11)))=(10(sqrt(21)-sqrt(11)))/((sqrt(21))^2-(sqrt(11))^2)=(10(sqrt(21)-sqrt(11)))/(21-11)=(10(sqrt(21)-sqrt(11)))/10=sqrt(21)-sqrt(11)

Power up,
HyperBrain!